OpenStudy (anonymous):

2 over -5+(68+14)+6^2

OpenStudy (anonymous):

do 6^2 then add that to 58=14 then simplify

OpenStudy (anonymous):

Im confused.. :(

OpenStudy (anonymous):

what is 6 squared

OpenStudy (anonymous):

36

OpenStudy (anonymous):

68+14=

OpenStudy (anonymous):

82. but what did you do with the fraction?

OpenStudy (anonymous):

we are adding the bottum together first

OpenStudy (anonymous):

so -5+82+36=

OpenStudy (anonymous):

113

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

no 113 is not the answer you still have to remeber the numerator of 2

OpenStudy (anonymous):

-2/5 + (68 + 14) + 6^2 Remember order of operations...PEMDAS do parenthesis first..... -2/5 + 82 + 6^2 next do exponents... -2/5 + 82 + 36 now add... -2/5 + 118 find common denominator which is 5 - 2/5 + 590/5 = 589/5 = 117 4/5